Please help me solve for X.
1.2x-7=-1.25x+8.6

Answer by LinnW(1048)

add 1.25x to each side
1.2x + 1.25x -7 = -1.25x + 1.25x +8.6
2.45x -7 = 8.6
add 7 to each side
2.45x = 15.6
divide each side by 2.45
x = 15.6/2.45
x = 6.36735